# Minimal Fractional Topological Insulator in half-filled conjugate moiré Chern bands

### v2

@@ Appendix A Classifying minimal TOs with or without time-reversal symmetry
In this appendix, we perform a systematic classification of minimal TOs for gapped topological states of electrons with fractional spin-Hall conductivity $\sigma^{\rm sh}=\frac{1}{q}$ (with positive integer $q$), but no diagonal charge or spin quantum Hall effect (i.e. no transverse spin current responding to spin gauge field). This classification of minimal TOs only assumes the charge and the spin-$S_{z}$ conservations. Then, we identify the mFTIs by further imposing the time-reversal symmetry ${\cal T}$. It turns out that there is a unique mFTI for each $q$. We separately discuss the cases with even and odd $q$’s. For even $q$, we find that there are exactly eight minimal TOs, all with squared total quantum dimension $\mathcal{D}^{2}=8q^{2}$. One of them is the $\mathds{Z}_{2q}$ TO, the same TO as the $\mathds{Z}_{2q}$ toric code. Further imposing time-reversal symmetry, we find that this $\mathds{Z}_{2q}$ TO is singled out as the unique mFTI that satisfies all the symmetry and transport requirements. For odd $q$, there is a unique minimal TO given by a product TO $\mathrm{U}(1)_{q}\boxtimes\mathrm{U}(1)_{-q}$, which can also be compatible with time-reversal symmetry.
@@ A.1 Minimal TOs without assuming time-reversal symmetry
Here, Spin$(n)_{1}$ denotes the TO of the level-1 Spin$(n)$ Chern-Simons theory, $\psi$ is the neutral fermion in Spin$(n)_{1}$, and $f$ represents the physical electron/hole. ${\bf e}$ and ${\bf m}$ are the two elementary anyons in the $\mathds{Z}_{2q}$ toric code $\mathrm{D}(\mathds{Z}_{2q})$. The quotient means condensing the bound state of ${\bf e}^{q}{\bf m}^{q}$, $\psi$, and $f$.
One can check that all the eight theories above are distinct from each other. Without further conditions, we have found all possible minimal TOs. Notice that even (odd) $n$’s corresponds to the type-I (type-II) minimal TOs. For example, the TO discussed in Sec. V.3 corresponds to $\mathcal{M}_{1}$.
@@ A.2 Imposing time-reversal symmetry
For the rest of the discussion, we focus on even $q$. Among ${\cal M}_{n=0,1,...,7}$ in Eq. (27), the odd $n$ theories all have non-zero chiral central charge and exhibit thermal Hall effect, breaking the time-reversal symmetry. So let us consider the even $n$ theories. The theory ${\cal M}_{0}$ is the mFTI discussed in Sec. II and Sec. V.4. As we will see in the following, any other theories ${\cal M}_{n\neq 0}$ cannot be made compatible with time-reversal symmetry.
For even $n$, $\mathcal{M}_{n}$ is an Abelian TO and can be described by a $K$-matrix theory. First, recall that the $\mathrm{U}(1)_{q}\boxtimes\mathrm{U}(1)_{-q}$ theory is described by the following Chern-Simons theory:

### v3

@@ Appendix A Classifying minimal TOs with or without time-reversal symmetry
In this appendix, we perform a systematic classification of minimal TOs for gapped topological states of electrons with fractional spin-Hall conductance $\sigma^{\rm sh}=\frac{1}{q}$ (with positive integer $q$), but no diagonal charge or spin quantum Hall effect (i.e. no transverse spin current responding to spin gauge field). This classification of minimal TOs only assumes the charge and the spin-$S_{z}$ conservations. Then, we identify the mFTIs by further imposing the time-reversal symmetry ${\cal T}$. It turns out that there is a unique mFTI for each $q$. We separately discuss the cases with even and odd $q$’s. For even $q$, we find that there are exactly eight minimal TOs, all with squared total quantum dimension $\mathcal{D}^{2}=8q^{2}$. One of them is the $\mathds{Z}_{2q}$ TO, the same TO as the $\mathds{Z}_{2q}$ toric code. Further imposing time-reversal symmetry, we find that this $\mathds{Z}_{2q}$ TO is singled out as the unique mFTI that satisfies all the symmetry and transport requirements. For odd $q$, there is a unique minimal TO given by a product TO $\mathrm{U}(1)_{q}\boxtimes\mathrm{U}(1)_{-q}$, which can also be compatible with time-reversal symmetry.
@@ A.1 Minimal TOs without assuming time-reversal symmetry
Here, Spin$(n)_{1}$ denotes the TO of the level-1 Spin$(n)$ Chern-Simons theory, $\psi$ is the neutral fermion in Spin$(n)_{1}$, and $f$ represents the physical electron/hole. ${\bf e}$ and ${\bf m}$ are the two elementary anyons in the $\mathds{Z}_{2q}$ toric code $\mathrm{D}(\mathds{Z}_{2q})$. The quotient means condensing the bound state of ${\bf e}^{q}{\bf m}^{q}$, $\psi$, and $f$. The reason why we only have $n=0,1,2,..,7$ is that ${\cal M}_{n}$ and ${\cal M}_{n+8}$ share the same anyon content up to a relabeling of anyon types.
One can check that all the eight theories above are distinct from each other. Without further conditions, we have found all possible minimal TOs. All the minimal TOs share the same minimal electric charge $e^{*}=e/q$. Notice that even (odd) $n$’s correspond to the type-I (type-II) minimal TOs. For example, the non-Abelian TO discussed in Sec. V.3 corresponds to $\mathcal{M}_{1}$ for $q=2$.
In the case of $q=2$, the type-I minimal TOs ${\cal M}_{0,2,4,6}$ and the type-II minimal TOs ${\cal M}_{1,3,5,7}$ follow the braiding statistics listed in Sec. IV. The extra data of the topological spin of the anyon $y$ can be obtained from Eq. (28):
$\theta_{y}=e^{\mathrm{i}n\pi/8}.$ (29)
Also, recall that $y$ is an Abelian anyon in ${\cal M}_{0,2,4,6}$ and a non-Abelian anyon in ${\cal M}_{1,3,5,7}$. From the value of $\theta_{y}$ and the fractional charge $q_{y}=(1/4,1/4)$, we can obtain the following fusion rule:
$y\times y=\begin{cases}v_{1}v_{2}^{3},&n=0,4,\\ v_{1}v_{2}f,&n=2,6,\\ v_{1}v_{2}^{3}+v_{1}v_{2}f,&n=1,3,5,7.\end{cases}$ (30)
The chiral central charge of these minimal TOs requires more careful treatment. Ref. Lapa and Levin, 2019 showed that, in a charge-conserved TO built only from charge-$e$ fermions, the chiral central charge $c$ can be determined by the anyon data of the TO and its charge fractionalization modulo 8. While ${\cal M}_{n}$ and ${\cal M}_{n+8}$ share the same anyon data, their charge fractionalization can still be different. A manifestation of such difference is that the topological spin $\theta_{y}$ in Eq. (29) of the anyon $y$ with fractional charge $(1/4,1/4)$ acquires an extra $-1$ factor when we change $n$ to $n+8$. If one further attaches an $f$ to $y$ to restore the topological spin of $y$, the fractional charge of $y$ is then altered. Hence, at the level of the topological order alone, the minimal TOs are classified by ${\cal M}_{n}$ with $n=0,1,...,7$. As minimal TOs with $\mathrm{U}(1)_{\uparrow}\times\mathrm{U}(1)_{\downarrow}$ symmetry fractionalization, we should allow the range of $n$ to be $n=0,1,2,...,15$. These 16 TOs can distinguished by $\theta_{y}$ (though the fusion rule of $y$ still only depends on $n$ mod 8 via Eq. (30)), with $q_{y}=(1/4,1/4)$. The chiral central charge of these TOs ${\cal M}_{n=0,1,...,15}$ are given by
$c=n/2~{}~{}{\rm mod}~{}8.$ (31)
Here the mod 8 chiral central charge is because one can stack a $E_{8}$ state formed out of neutral bosons, without changing the bulk SET order.
@@ A.2 Imposing time-reversal symmetry
For the rest of the discussion, we focus on even $q$. From the chiral central charge Eq. (31), only $n=0$ is compatible with time-reversal symmetry. The theory ${\cal M}_{0}$ is the mFTI discussed in Sec. II and Sec. V.4. Hence, the mFTIs with $\mathbb{Z}_{2q}$ TOs are the unique mFTIs.
For even $n$, we can provide a different analysis that also confirms that ${\cal M}_{0}$ is the only minimal TO compatible with time-reversal symmetry. With $n$ even, $\mathcal{M}_{n}$ is an Abelian TO and can be described by a $K$-matrix theory. First, recall that the $\mathrm{U}(1)_{q}\boxtimes\mathrm{U}(1)_{-q}$ theory is described by the following Chern-Simons theory: